OpenStudy (anonymous):

Solve for the roots in the equation below. x^4 + 3x^2 - 4 = 0

OpenStudy (anonymous):

x=2i,-2i,1,-1

OpenStudy (anonymous):

Can you show how you got x=2i,-2i,1,-1? @GotToLoveJenny

OpenStudy (anonymous):

(x^2 +4 ) *(x^2 -1 ) =0 x^2 -1 =0 (x-1)(x+1)=0 x-1=0 so x=1 x+1=0 so x=-1 x^2 +4=0 x^2=-4 \[x=\sqrt{-4}\] it is not found as real but as imaginary x=+2i ,x=-2i as i=\[\sqrt{-1}\]

OpenStudy (anonymous):

\[\sqrt{-4}=\sqrt{4}*\sqrt{-1}=2*\sqrt{-1}\]

OpenStudy (anonymous):

thanks hanishala

OpenStudy (anonymous):

you are welcome @GotToLoveJenny

OpenStudy (anonymous):

so the roots are?

OpenStudy (anonymous):

2i ,-2i ,1 ,-1

OpenStudy (anonymous):

the roots are sqrt-4,+2i,-2i, sqrt-1?

OpenStudy (anonymous):

or 2i, -2i, 1, -1?

OpenStudy (anonymous):

2i ,-2i ,-1 ,1 @Power2Knowledge

OpenStudy (anonymous):

Thank you, my pc had glitched.